OpenStudy (anonymous):

P can do 1/4 of a work in 10 days, Q can do 40% of the work in 15 days, and R can do 1/3 of the work in 13 days. Who will complete the work first? Please show workings.

OpenStudy (anonymous):

Q will complete the work first in 37.5 days.

OpenStudy (anonymous):

can you show the workings?

OpenStudy (anonymous):

40/100 = 2/5 2/5 of x = 15 x is total work x = 15*5/2 = 75/2 = 37.5 days For P and R also in same way and P can do the work in 40 days and R in 39 days.

OpenStudy (anonymous):

thanks.

OpenStudy (anonymous):

Welcome